Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

1. Section 6.4: problem 5.

2. Chap 6, Additional and Advanced Exercises: problems 4.

3. Section 7.2: Problems 13, 19, 21.

4. Section 9.2: problems 3, 9, 19, 31 (optional).